OpenStudy (loser66):

I don't see how: if \(r_1(t) = (1-t)P+t (1,0) ~~~0\leq t\leq1\) \(r_2(t) =(1-t)(1,0) + tQ~~~0\leq t\leq 1\) where P, Q are initial and terminal points respectively, then \((r_1+r_2)(t)=\begin {cases}r_1(2t) ~~~0\leq t\leq 1/2\\r_2(2t-1)~~1/2\leq t\leq 1\end {cases}\) Please, help

OpenStudy (loser66):

@dan815

OpenStudy (loser66):

@freckles

OpenStudy (freckles):

if we have r(t)=f(t) on the interval 0<t<1 then r(2x)=f(2x) on the interval 0<2x<1 so that would mean 0<x<1/2 or if we had r(2x-1)=f(2x-1) the 0<2x-1<1 so 1<2x<2 then 1/2<x<1

OpenStudy (freckles):

is that what you are asking about the intervals?

OpenStudy (freckles):

or the whole thing

OpenStudy (loser66):

the whole. :)

OpenStudy (freckles):

Well I understand how they got the intervals but the other parts...

OpenStudy (freckles):

\[(r_1+r_2)(t)=(1-t)P+t(1,0)+(1-t)(1,0)+tQ\]

OpenStudy (loser66):

=P -tP +(1,0) +tQ :)

OpenStudy (freckles):

\[r_1(2t)=(1-2t)P+2t(1,0)\]

OpenStudy (loser66):

but the interval is changed.

OpenStudy (freckles):

\[r_2(2t-1)=(1-(2t-1))(1,0)+(2t-1)Q \\ = (-2t+2)(1,0)+(2t-1)Q\]

OpenStudy (freckles):

and yeah so the interval is changed ...

OpenStudy (freckles):

stupid question (1,0) means what exactly

OpenStudy (xapproachesinfinity):

hmmm interesting

OpenStudy (freckles):

oh wait

OpenStudy (freckles):

(1,0) is a point and Q is a point

OpenStudy (freckles):

so r_2 is a point right?

OpenStudy (freckles):

r_1 is a point

OpenStudy (loser66):

Those are from my Prof's lecture. I don't get this part. :)

OpenStudy (freckles):

it looks like they just translated r_1 and r_2

OpenStudy (freckles):

\[r_1(2t)=(1-2t)P+2t(1,0) , 0 \le t \le \frac{1}{2} \equiv r_1(t)=(1-t)P+t(1,0),0\le t \le 1\] (1-t) is the same as (1-2x) when t=0 and x=0 (1-t) is the same (1-2x) when t=1 and x=1/2

OpenStudy (loser66):

this is original problem:

OpenStudy (freckles):

I plugged in 1/4 one way I got \[(1,0)+\frac{3}{4}P+\frac{1}{4}Q \\ \] and the other way I got \[\frac{1}{2}P+\frac{1}{2}(1,0)\] If i didn't make a sense. I don't know how to tell if these are the same :(

OpenStudy (freckles):

if i didn't make a mistake*

OpenStudy (freckles):

\[\frac{1}{4}Q=-\frac{1}{2}(1,0)-\frac{1}{4}P \\ Q=-2(1,0)-P?\]

ganeshie8 (ganeshie8):

aren't you just changing the parameter so that t = (0,1) gives you two sequential paths : 0->1/2 : r1 1/2 -> 1: r2

OpenStudy (loser66):

Please, explain more.@ganeshie8

ganeshie8 (ganeshie8):

if you think of "t" as time and both the paths as the walk of bugs : 

ganeshie8 (ganeshie8):

the first parameterization gives you two parallel paths : two bugs starting at different positions and travelling as described by the path

ganeshie8 (ganeshie8):